### v1

# Informed Initialization for Bayesian Optimization and Active Learning

Abstract
Bayesian Optimization is a widely used method for optimizing expensive black-box functions, relying on probabilistic surrogate models such as Gaussian Processes. The quality of the surrogate model is crucial for good optimization performance, especially in the few-shot setting where only a small number of batches of points can be evaluated. In this setting, the initialization plays a critical role in shaping the surrogate’s predictive quality and guiding subsequent optimization. Despite this, practitioners typically rely on (quasi-)random designs to cover the input space. However, such approaches neglect two key factors: (a) space-filling designs may not be desirable to reduce predictive uncertainty, and (b) efficient hyperparameter learning during initialization is essential for high-quality prediction, which may conflict with space-filling designs. To address these limitations, we propose Hyperparameter-Informed Predictive Exploration (HIPE), a novel acquisition strategy that balances predictive uncertainty reduction with hyperparameter learning using information-theoretic principles. We derive a closed-form expression for HIPE in the Gaussian Process setting and demonstrate its effectiveness through extensive experiments in active learning and few-shot BO. Our results show that HIPE outperforms standard initialization strategies in terms of predictive accuracy, hyperparameter identification, and subsequent optimization performance, particularly in large-batch, few-shot settings relevant to many real-world Bayesian Optimization applications.
1 Introduction
Bayesian Optimization (BO) (Mockus et al., 1978; Jones et al., 1998; Frazier, 2018; Garnett, 2023) is a principled framework for sample-efficient global optimization of black-box functions with applications across diverse fields such as biological discovery (Griffiths and Hernández-Lobato, 2020; Stanton et al., 2022), materials science (Frazier and Wang, 2016; Attia et al., 2020; Ament et al., 2023b), online A/B testing (Agarwal et al., 2018; Letham et al., 2019; Feng et al., 2025), and machine learning hyperparameter optimization (HPO) (Snoek et al., 2014; Feurer et al., 2015). BO combines a probabilistic surrogate model—commonly a Gaussian Process (GP)—with an acquisition function to select where to evaluate the unknown objective function. In many applications, the runtime of a single black-box function evaluation may restrict the experimenter to a small number of batches – the number of sequential rounds of experiments – but many real-world experimental setups permit conducting multiple experiment simultaneously (e.g., on a parallel compute cluster, in a randomized controlled trial, or batch-testing multiple specimens in a lab).
The success of Bayesian Optimization in practice is highly sensitive to the quality of the surrogate model (Eriksson and Jankowiak, 2021; Hvarfner et al., 2023; Foldager et al., 2023). This is a challenge particularly during the early stages of optimization when few observations are available. In these early stages, a small set of inputs is typically selected at random, or via space-filling or quasi-random sampling strategies such as Latin Hypercube Sampling (LHS) or scrambled Sobol’ sequences (Bossek et al., 2020; Owen, 2023). While such strategies aim to achieve broad coverage of the input space for initialization of the surrogate model, they may not necessarily result in improving predictive accuracy, or reduce predictive variance, across the entire input space Ren and Sweet (2024). Moreover, they neglect a a second crucial aspect of modeling: the need to accurately infer the model’s hyperparameters (Zhang et al., 2019), such as the kernel lengthscales of a GP. With accurate hyperparameter estimates, variation in unimportant dimensions will have less influence on the selection of points in subsequent iterations of BO, leading to more sample-efficient optimization (Eriksson and Jankowiak, 2021; Hvarfner et al., 2023; Müller et al., 2023). Conversely, poor hyperparameter estimation may cause subsequent BO iterations to fail to make meaningful progress (Berkenkamp et al., 2019), which may be caused by misidentifying signal for noise, exploring irrelevant dimensions, or returning poor terminal recommendations (Hvarfner et al., 2022).
In this work, we provide a principled approach to addressing this initialization challenge. Our main contributions are as follows:
  1. 1.
    We propose Hyperparameter-Informed Predictive Exploration (HIPE), an acquisition function for initialization that optimizes for both predictive uncertainty reduction and hyperparameter learning.
  2. 2.
    We derive a closed-form expression for this objective in the case of Gaussian Process models, and implement a practical Monte Carlo approximation to make it amenable to batched optimization.
  3. 3.
    We conduct extensive experiments in active learning and Bayesian Optimization on synthetic and real-world BO tasks, demonstrating that HIPE outperforms competing methods in terms of both model accuracy metrics and BO performance in few-shot, large-batch settings.
2 Background
2.1 Gaussian Processes
Gaussian Processes (GPs) are a widely used surrogate model in BO due to their flexibility, closed-form and well-calibrated predictive distributions. GPs define a distribution over functions, $\hat{f}\sim\mathcal{GP}(m(\cdot),k(\cdot,\cdot))$, specified by a mean function $m(\cdot)$ and a covariance (kernel) function $k(\cdot,\cdot)$. For a given location $\bm{x}$, the function value $\hat{f}(\bm{x})$ is normally distributed, with closed-form expressions for the predictive mean $\mu(\bm{x})$ and variance $\sigma^{2}(\bm{x})$. In practice, the mean function is often kept constant, leaving the covariance function to capture the structural properties of the objective.
To model differences in variable importance in GPs with stationary kernels, each input dimension is commonly scaled by a lengthscale hyperparameter $\ell_{i}$, a practice known as Automatic Relevance Determination (ARD) (Williams and Rasmussen, 1995). Additional, optional hyperparameters include a learnable noise variance $\sigma_{\varepsilon}^{2}$ and signal variance $\sigma_{f}^{2}$. The full set of hyperparameters $\bm{\theta}=\{\bm{\ell},\sigma_{\varepsilon}^{2},\sigma_{f}^{2}\}$ may be learned either by maximizing the marginal likelihood $p(\mathcal{D}\mid\bm{\theta})$ (Maximum Likelihood Estimation, MLE) or by incorporating hyperpriors $p(\bm{\theta})$ to perform Maximum A Posteriori (MAP) estimation. Alternatively, a fully Bayesian treatment (Osborne, 2010; Lalchand and Rasmussen, 2020) integrates over $\bm{\theta}$ to approximate the full Bayesian posterior distribution using Markov Chain Monte Carlo (MCMC) methods, thereby explicitly accounting for hyperparameter uncertainty. For additional background on GPs, see (Rasmussen and Williams, 2005).
2.2 Bayesian Optimization
Bayesian Optimization (BO) (Frazier, 2018) is a sample-efficient framework for finding the maximizer $\bm{x}_{*}=\mathop{\mathrm{arg\,max}}_{\bm{x}\in\mathcal{X}}f(\bm{x})$ of a black-box function $f:\mathcal{X}\rightarrow\mathbb{R}$ over a $D$-dimensional input space $\mathcal{X}=[0,1]^{D}$. The function $f$ is assumed to be expensive to evaluate and observable only through noisy point-wise measurements, $y(\bm{x})=f(\bm{x})+\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2})$.
At the core of BO is an acquisition function, which uses a surrogate model to quantify the (expected) utility of candidate points $\bm{x}$. Acquisition functions balance exploration and exploitation typically through greedy heuristics. Popular examples include Expected Improvement (EI) (Jones et al., 1998; Bull, 2011) and its numerically stable variant, LogEI (Ament et al., 2023a), as well as the Upper Confidence Bound (UCB) (Srinivas et al., 2012, 2010). In batch BO, multiple points are selected in parallel to accelerate data collection. This is often achieved by computing and jointly optimizing a (quasi-)MC estimate of the utility $u$ associated with acquisition function over the full batch ${\bm{X}}=\{\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{q}\}$ of size $q$ (Wilson et al., 2017, 2020; Balandat et al., 2020).
2.3 Bayesian Active Learning
Bayesian Active Learning (BAL) and Bayesian Experimental Design (BED) (Chaloner and Verdinelli, 1995) aims to improve predictive models by strategically selecting data points that are most informative, either with regard to the model or to future predictions. A central quantity is the Expected Information Gain (EIG):
$\textsc{EIG}(\xi;y(\bm{x}))=\text{H}[\xi]-\mathbb{E}_{y(\bm{x})}\left[\text{H}[\xi|y(\bm{x})]\right],$ (1)
where H is the Shannon (differential) entropy and $\xi$ is a parameter of interest. Importantly, EIG is symmetric, and can be equivalently formulated as an entropy reduction over $y(\bm{x})$, instead of $\xi$.
Bayesian Active Learning by Disagreement
Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011; Kirsch et al., 2019) selects query points that maximize the mutual information between model predictions and hyperparameters $\bm{\theta}$:
$\text{BALD}(\bm{x})=\textsc{EIG}(y(\bm{x});\bm{\theta})=\text{H}[y(\bm{x})|\mathcal{D}]-\mathbb{E}_{\bm{\theta}}[\text{H}[y(\bm{x})|\mathcal{D},\bm{\theta}]].$ (2)
BALD identifies locations where models within an ensemble exhibit the greatest disagreement in predictive uncertainty. In the GP setting, this often leads to axis-aligned queries when there is high uncertainty in the lengthscales, and to repeated queries when observation noise is highly uncertain. Notably, BALD is model-agnostic and can be applied to a wide range of surrogate models and hyperparameters, including subspace models (Garnett et al., 2014) and additive decompositions (Hvarfner et al., 2023; Gardner et al., 2017).
Negative Integrated Posterior Variance and Expected Predictive Information Gain
The Negative Integrated Posterior Variance (NIPV) (Chen and Zhou, 2014; Seo et al., 2000) criterion selects queries that minimize the expected posterior variance over a test distribution $p_{*}(\bm{x})$:
$\text{NIPV}(\bm{x};p_{*})=-\mathbb{E}_{\bm{x}_{*}\sim p_{*}}\left[\sigma^{2}(\bm{x}_{*})\;|\ \bm{x},\mathcal{D}\right].$ (3)
Similarly, Expected Predictive Information Gain (EPIG) (Bickford Smith et al., 2023) selects queries that minimize the expected predictive entropy over $p_{*}(\bm{x})$:
$\text{EPIG}(\bm{x};p_{*})=-\mathbb{E}_{\bm{x}_{*}\sim p_{*}}\left[\text{H}[y(\bm{x})]\;|\;\bm{x}_{*},\mathcal{D}\right].$ (4)
Both objectives promote the selection of data that reduces uncertainty over the test distribution, but without considering the effect on hyperparameter learning. The test distribution encodes how much emphasis is put on different parts of the domain. It can be specified by subject matter experts; in the case of no prior knowledge it is typically the uniform distribution. Throughout the remainder of the paper, we will exclusively consider NIPV with a uniform $p_{*}$.
3 Related Work
Initialization has received surprisingly little attention in the context of BO. Bossek et al. (2020) conducts a study on the effect of various random initial designs on BO performance. Alternatively, minimax or maximin criteria (Johnson et al., 1990) may be used to accomplish evenly distributed designs. Maybe closest to our work is (Zhang et al., 2019), which proposes LHS-Beta, an initial design criterion which alters samples drawn by LHS to achieve pairwise distances between points which matches a Beta distribution. LHS-Beta pursues diverse pairwise distances in the data, in order to best learn the lengthscale of a GP with an isotropic kernel. Zimmerman (2006); Müller and Zimmerman (1999) address the problem of learning parameters of Kriging estimators using a empirical estimates of optimal experimental design criteria, limiting candidates to a fixed grid of points.
Entropy-maximizing (Guestrin et al., 2005; MacKay, 1992, 1995) or variance-minimizing (Park et al., 2024) designs have been explored in active learning for optimal sensor placement (Krause et al., 2006, 2008) and other applications involving GPs, such as geostatistics (Sauer et al., 2023) and contour finding (Cole et al., 2023). Moreover, parameter-related EIG criteria are a bedrock of the broader topic of BED (Bickford Smith et al., 2023; Rainforth et al., 2024; Kirsch et al., 2021; Chaloner and Verdinelli, 1995), which focuses on selecting data that is most informative about model parameters or future predictions. These prediction or model-oriented criteria have yet to see widespread use in BO, particularly for initialization. However, information-theoretic acquisition functions (Hennig and Schuler, 2012; Hernández-Lobato et al., 2014; Wang and Jegelka, 2017; Moss et al., 2021; Hvarfner et al., 2022; Tu et al., 2022; Neiswanger et al., 2021, 2024) address the optimization problem from an information theoretic perspective, albeit not with a primary focus on initialization or model predictive performance.
The problem of actively learning model hyperparameters during BO (post initialization) has previously been investigated by Hvarfner et al. (2023), who propose a combined BO-BAL framework to actively learn the hyperparameters of the GP along with the optimum, and demonstrate that better-calibrated surrogates significantly enhance BO performance. Houlsby et al. (2011) proposes BALD, an active learning acquisition functions for hyperparameters in preference learning in GPs. Riis et al. (2022) proposes a a Query-By-Committee-oriented acquisition function for BAL in GPs. Lastly, Berkenkamp et al. (2019); Ziomek et al. (2024) address BO performance under hyperparameter uncertainty from a theoretical perspective, proving regret bounds when hyperparameters of the objective are unknown.
4 Method
We consider the case of large-batch, few-shot BO, where the batch size $q$ is large ($q\geq 8$), and only a small number of sequential batches $B$ can be evaluated (often only one for initialization and one for BO, so $B=2$). With only a handful of batches, there is little opportunity to correct for poor initialization. If the initial design fails to reduce uncertainty in key hyperparameters or leaves large regions of the input space unexplored, both hyperparameter inference and optimization become challenging. This setting is common if evaluations are lengthy but can be effectively parallelized, which is the case for instance in online A/B tests or lab experiments for biology Stanton et al. (2022); González-Duque et al. (2024) or material design Yik et al. (2025).
[…]
While Sobol and other space-filling sequences are designed to uniformly cover the input space, this property alone does not imply that the resulting model will be desirably helpful for subsequent BO. In fact, space-filling is not synonymous with informativeness: a model trained on a space-filling design may still exhibit high predictive uncertainty in regions most relevant for optimization. For example, random designs increasingly sample near the boundaries of the search space as dimensionality grows Swersky (2017); Köppen (2000), whereas a point located at the center of the search space may be far more informative for optimization, as it can yield a greater reduction in predictive variance throughout the search space.
To this point, a strategy that explicitly reduces predictive uncertainty within the region of interest Bickford Smith et al. (2023)—namely, the search space—not only seems better suited for our goals, but is also better aligned with both theoretical (Srinivas et al., 2010; Berkenkamp et al., 2019) and information gain-based criteria, as well as the practical objective of achieving a more informed, in an informal sense, model post-initialization Garnett (2023). Methods like NIPV (Zhang et al., 2019) which minimize predictive variance over the input space achieve this goal better than Sobol, and yields designs that are spread out but not necessarily informative for model calibration and hyperparameter learning.
On the other hand, methods that focus exclusively on hyperparameter informativeness, such as BALD (Houlsby et al., 2011), tend to cluster queries along specific axes in order to resolve lengthscales, thereby sacrificing coverage and potentially overlooking important regions of the search space. When additional hyperparameters are present, such as a learnable noise variance, BALD may even select duplicate queries at the same location to better estimate said noise variance. While such behavior is not inherently problematic, it can be undesirable in the few-shot setting, where initialization must address hyperparameter learning and space-filling simultaneously.
To overcome these limitations, we introduce HIPE, a method that explicitly balances predictive uncertainty reduction with hyperparameter-awareness. By jointly considering coverage and informativeness, HIPE produces initial designs that effectively reduce both predictive and hyperparameter uncertainty, ensuring robust model calibration and improved downstream optimization.
The interplay between all these initialization strategies is illustrated in Figure 1, which visualizes the acquisition surfaces for a GP model in two dimensions under lengthscale uncertainty. Each subplot highlights the distinct behavior of a different method. Sobol emphasizes uniform coverage of the input space, but its designs may not achieve the desired reduction in predictive uncertainty, especially given the inherent randomness and lack of model awareness. BALD, in contrast, focuses on reducing lengthscale uncertainty by selecting axis-aligned queries, which can lead to clustering along specific directions and a loss of broader coverage. NIPV aims to minimize average predictive variance across the input space, resulting in well-spread points, but it can neglect hyperparameter informativeness since its queries are not tailored to resolve model parameter uncertainty. Our proposed method, HIPE, strikes a balance between predictive uncertainty reduction and hyperparameter-awareness: it selects points that are both well-distributed and aligned with the axes of uncertainty, thereby achieving low predictive and hyperparameter uncertainty simultaneously.
Figure 1: Visual comparison of initialization strategies for BO with a GP in two dimensions under lengthscale uncertainty. Each plot shows the acquisition surface used to guide batch selection. Sobol emphasizes space-filling, but may not accomplish this to the desired degree. BALD focuses on reducing lengthscale uncertainty via axis-aligned queries. NIPV spreads points to minimize average predictive variance over the input space. Lastly, HIPE balances space-filling and hyperparameter-awareness, choosing spread-out points while preserving axis-alignment between queries.
4.1 Hyperparameter-Informed Predictive Exploration
We approach the problem of initialization through the lens of optimization, by simultaneously maximizing the coverage of a region of interest and the information acquired about model-level uncertainty. A natural way to achieve this is to optimize a criterion that combines the EPIG and BALD objectives:
$\text{{HIPE}}_{\beta}({\bm{X}}):=\underbrace{-\mathbb{E}_{\bm{\theta},y({\bm{X}})}\left[\mathbb{E}_{\bm{x}_{*}}\left[\text{H}\left[y(\bm{x}_{*})\mid\bm{\theta},y({\bm{X}})\right]\right]\right]}_{\text{EPIG objective}}+\beta\;\underbrace{\left(\text{H}[y({\bm{X}})]-\mathbb{E}_{\bm{\theta}}\left[\text{H}[y({\bm{X}})\mid\bm{\theta}]\right]\right)}_{\text{BALD objective}}$ (5)
where $\beta>0$ is a scalar weighting parameter. For large $\beta$, this objective favors hyperparameter learning, while for small $\beta$ it favors space-filling designs.
It turns out that for the choice of $\beta=1$, the maximizer of Eq. (7) is exactly the maximizer of the joint information gain over test function values and model hyperparameters (for proof see Appendix B):
Proposition 1 (Equivalence of $\textsc{HIPE}_{\beta=1}$ to Joint Information Gain).
The $\textsc{HIPE}_{\beta}$ acquisition function with $\beta=1$ is equivalent to maximizing the expected joint information gain over test function values $y(\bm{x}*)$ and model hyperparameters $\bm{\theta}$ acquired by a candidate batch $\bm{X}$}. Formally,
$\mathop{\mathrm{arg\,max}}_{{\bm{X}}\in\mathbb{R}^{q\times D}}\;\textsc{HIPE}_{\beta=1}({\bm{X}};p_{*})=\mathop{\mathrm{arg\,max}}_{{\bm{X}}\in\mathbb{R}^{q\times D}}\;\mathbb{E}_{\bm{x}\sim p_{*}}\left[\textsc{EIG}\left(y(\bm{x}_{*}),\bm{\theta};{\bm{X}}\right)\right].$ (6)
While Proposition 1 provides an intuitive connection between HIPE, EPIG, and BALD, the constituent quantities often have vastly different scales. Therefore the choice of $\beta=1$ will generally not result in optimal performance since the optimization will inadvertently focus on the larger of the two terms.
A key observation is that not all hyperparameter information gain amounts to information gained on the test set – this depends on multiple aspects, including downstream test distribution and hyper-parameterization. To what extent the reduction in hyperparameter entropy manifests in a reduction in test set entropy can be quantified through the mutual information between the hyperparameters $\bm{\theta}$ and the test points $y(\bm{x}_{*}),\;\bm{x}_{*}\sim p_{*}(\bm{x})$:
$\textsc{EIG}(y(\bm{x}_{*});\bm{\theta}|\mathcal{D})=\mathbb{E}_{\bm{x}_{*}}\left[\text{H}[y(\bm{x}_{*})|\mathcal{D}]-\mathbb{E}_{\bm{\theta}}[\text{H}[y(\bm{x}_{*})|\bm{\theta},\mathcal{D}]]\right].$ (7)
Intuitively, Eq. (7) quantifies how well the knowledge of the hyperparameters, in expectation, informs us about the values of $y(\bm{x}_{*})$. Importantly, $\textsc{EIG}(y(\bm{x}_{*});\bm{\theta}|\mathcal{D})$ does not depend on the candidate set $\bm{X}$} and can thus be pre-computed.
Setting $\beta=\textsc{EIG}(y(\bm{x}_{*});\bm{\theta}|\mathcal{D})$ balances the two competing objectives in Eq. (5) according to their effect on downstream predictive uncertainty, without introducing any additional hyperparameters. We refer to the resulting acquisition function as Hyperparameter-Informed Predictive Exploration (HIPE):
$\text{{HIPE}}({\bm{X}}):={-\mathbb{E}_{y({{\bm{X}}})}\left[\mathbb{E}_{\bm{x}_{*}}[\text{H}[y(\bm{x}_{*})|\bm{\theta},y({{\bm{X}}})]\right]}+\textsc{EIG}(y(\bm{x}_{*});\bm{\theta})\;{\mathbb{E}_{\bm{\theta}}[\text{H}[y({{\bm{X}}})|\bm{\theta}]-\text{H}[y({{\bm{X}}})]]}$ (8)
Notably, optimizing HIPE does not require having observed any data – for this problem to be well-posed only requires test distribution, model structure, and model parameter hyperpriors.
4.2 A Parallel Monte Carlo Implementation of HIPE
Following the Monte Carlo approach used in modern acquisition functions (Wilson et al., 2020; Balandat et al., 2020), we implement a parallel version of HIPE that enables joint optimization. For a candidate batch ${\bm{X}}=\{\bm{x}_{1},\dots,\bm{x}_{q}\}$, we estimate the acquisition objective using $M$ MC samples over the hyperparameters, $T$ test locations of the EPIG objective, and $N$ samples from the predictive posterior for the BALD objective:
$\alpha_{\text{BALD}}({\bm{X}};\bm{\theta})$ $=-\log p\left(\!\mathbb{E}_{\bm{\theta}}[y({\bm{X}})|\bm{\theta})]\!\right)+\mathbb{E}_{\bm{\theta}}\left[\log p(y({\bm{X}})\mid\bm{\theta})\right]$ (9a)
$\approx-\frac{1}{N}\sum_{n=1}^{N}\!\left[\log\left(\frac{1}{M}\!\sum_{m=1}^{M}p\left(Y^{(n)}\mid\theta^{(m)}\right)\right)+\frac{1}{M}\!\sum_{m=1}^{M}\log p\left(\!Y^{(n)}\mid\theta^{(m)}\!\right)\right],$ (9b)
where $\bm{\theta}^{(m)}\sim p(\bm{\theta}\mid\mathcal{D})$ are i.i.d samples from the belief over hyperparameters, and $Y^{(n)}\sim y({\bm{X}})$ are i.i.d, $q$-dimensional (joint) samples from the predictive posterior. Thus, the BALD objective amounts to repeated evaluation the $q$-dimensional multivariate normal posterior $y({\bm{X}})$ for each candidate $\bm{X}$}, and estimating the posterior entropy from it. Secondly, we estimate the EPIG objective analogously to Balandat et al. (2020) as
$\alpha_{\text{EPIG}}({\bm{X}};p_{*},\bm{\theta})\approx\frac{1}{M}\frac{1}{T}\sum_{m=1}^{M}\sum_{t=1}^{T}\left[C-\text{H}[y(\bm{x}_{*}^{(t)})|{\bm{X}},\bm{\theta}^{(m)},\mathcal{D}]\right],$ (10)
where $C=\text{H}[y(\bm{x}_{*}^{(t)})|\mathcal{D}]$ is constant w.r.t. $\bm{X}$} and does not need to be computed. Since $y(\bm{x}_{*}^{(t)})|{\bm{X}},\bm{\theta}^{(m)},\mathcal{D}$ in Eq. (10) is the GP posterior predictive at $\bm{x}_{*}$, it is a Gaussian random variable and its entropy can be computed in closed form without observing or simulating the outcome at $\bm{x}_{*}$. Thus, Eq. (10) can be evaluated in closed form for each of the $M$ GPs’ $T$ test points after conditioning on the set $\bm{X}$}.
With these two MC estimators, the subsequent optimization can be carried out jointly for the entire $q$-batch in a $qD$-dimensional space, as the entropy reduction is computed with regard to the entire batch of candidates as opposed to a singular one. Using Sample Average Approximation (Balandat et al., 2020), the HIPE objective is deterministic and auto-differentiable.
One downside of this formulation is that the nested MC estimator imposes substantial computational runtime, making HIPE less suited for high-throughput applications (Eriksson et al., 2019; Maus et al., 2022; Daulton et al., 2022). However, in those applications the quality of the initialization batch is generally much less crucial, so this is not a limitation in practice.
5 Results
We evaluate HIPE across two main types of tasks: Active Learning (AL) and Bayesian Optimization (BO). We consider various synthetic and real-world problems and a number of different baselines. In both settings, we maintain large batch sizes and few batches.
Setup
[…]
Evaluation Criteria
We measure model fit quality with Root Mean Square Error (RMSE) of the mean prediction and Negative Log-Likelihood (NLL) against a large number of (ground truth) test point sampled uniformly from the domain. In the “two-shot” optimization setting, we are interested in how the initialization affects the quality of the GP surrogate after both the first (initialization) and second (BO) batch. We compute relative rankings, the performance of each algorithm compared to its competitors, for each seed of each function, and average across the task type. As such, the relative rankings aggregate inter-algorithm performance across rows for Figs 24.
We also study how these model quality improvements translate to better optimization performance. To this end, we consider the out-of-sample inference performance (Hernández-Lobato et al., 2014; Hvarfner et al., 2022), that is, the performance of the point $\bm{x}^{\prime}=\mathop{\mathrm{arg\,max}}\mu(\bm{x}\mid\mathcal{D})$ selected as the maximizer of the posterior mean of the surrogate model fit on the data available in each batch. We choose this metric since using observed points directly performs very poorly in noisy settings, and only considering in-sample points is rather limiting in the few-shot setting.
5.1 Batch Active Learning on GP Surrogates and Synthetic Functions
We first evaluate the ability of HIPE to learn accurate surrogate models through batch active learning on noisy synthetic test functions and surrogate LCBench (Zimmer et al., 2021) tasks. The LCBench tasks are derived from complete neural network training runs on various OpenML (Vanschoren et al., 2014) datasets, with 7D GP surrogate models fitted as described in Appendix A. Additionally, we evaluate performance on the Hartmann 66D function and a high-dimensional Hartmann 66D (1212D) variant, where dummy input dimensions are added following standard practice in high-dimensional BO (Eriksson and Jankowiak, 2021). These dummy dimensions introduce an additional challenge, as effectively identifying and ignoring irrelevant features is critical for accurate predictions. All evaluations are subject to substantial observation noise, detailed in Appendix A.
We run each algorithm for $4$ batches of size $q=16$ and measure the NLL and RMSE after each batch, displaying mean and one standard error on all tasks. In Fig. 2 we see that, across all tasks, HIPE is the only method that consistently ranks in the top two for both RMSE and NLL, demonstrating that the models it produces are both accurate and well-calibrated. On NLL, HIPE performs comparably to BALD, which targets hyperparameter learning and thus excels at model calibration. Similarly, HIPE is competitive with NIPV on RMSE, a metric for which NIPV is particularly well-suited (Gramacy and Lee, 2009).
Figure 2: Model accuracy results in the batch active learning setting. We report RMSE across various synthetic and LCBench surrogate tasks over 4 batches of $q=16$ evaluations across 100100 seeds per benchmark. HIPE consistently ranks in the top two in relative rankings on both metrics, achieving a strong balance between hyperparameter learning and predictive accuracy. BALD performs competitively on marginal log-likelihood (MLL) but underperforms on RMSE due to limited space-filling behavior, while NIPV excels at reducing RMSE but struggles with model calibration. On aggregate, random initialization methods lag significantly behind across all benchmarks.
5.2 Noisy Synthetic Test Functions
Next, we evaluate HIPE on synthetic benchmark functions in the two-shot Bayesian optimization setting, using $B=2$ batches and a batch size of $q=24$. We consider three standard test functions—Ackley (44D), Hartmann (44D), and Hartmann (66D)—as well as two higher-dimensional variants of the Hartmann function. Observation noise is added to all tasks ($\sigma_{\varepsilon}=2$ for Ackley and $\sigma_{\varepsilon}=0.5$ for Hartmann), further increasing task difficulty. In Fig. 3 shows that across all benchmarks, HIPE consistently achieves the best or second-best performance, followed by NIPV. Random and space-filling initialization methods (LHS-Beta, Random, and Sobol) perform noticeably worse across all settings.
Figure 3: Out-of-sample inference optimization performance on noisy synthetic benchmark functions under the two-shot setting ($B=2$, $q=24$) across 100100 seeds per benchmark. HIPE outperforms or matches the best-performing method across all benchmarks, including on the high-dimensional Hartmann variants with added dummy variables. NIPV performs well on most tasks, but its performance degrades on tasks where hyperparameter identification is critical. Random initialization strategies perform the worst throughout.
5.3 LCBench HPO Tasks
We evaluate five additional tasks from LCBench in the two-shot optimization setting: Fashion-MNIST, MiniBooNE, Car, Higgs, and Segment, using fixed, minimal observation noise. Fig. 4 displays that HIPE substantially outperforms the competition on Fashion-MNIST, Mini-BooNE and Higgs, and performs competitively on the remaining Segment and Car. Overall, HIPE consistently delivers the highest relative rank, outperforming competing algorithms by a substantial margin. In App. 7, we demonstrate the performance of the same initialization schemes on $q=8$, $B=5$. In this setting, HIPE is the top-performing method on all benchmarks, and the overall ranking of algorithms remains roughly intact.
Figure 4: Two-shot optimization results on five hyperparameter optimization tasks from LCBench across 200 seeds for each task. HIPE achieves the highest final performance on four of the five tasks and remains competitive on the remaining one (Segment). Relative rankings across the five problems show that HIPE consistently outperforms other methods. This confirms the practical relevance of our approach for real-world HPO scenarios, where both accurate predictive modeling and effective exploration are crucial.
5.4 Runtime Analysis
We compare the time required to optimize HIPE to other options, namely NIPV, as well as its runtime compared to other components of a BO loop that employs fully Bayesian GP modeling. In Table 1, we demonstrate that, while HIPE is substantially slower to optimize than NIPV, its total runtime is still a minor component of the overall runtime of the BO run.
Runtime Component (sec) HIPE NIPV Random
Model fit (initial) 21.91 $\pm$ 0.15 21.91 $\pm$ 0.15 N/A
Acquisition Opt. 19.61 $\pm$ 1.12 8.62 $\pm$ 0.37
Model fit (BO loop) 181.79 $\pm$ 7.60 189.90 $\pm$ 5.17 177.33 $\pm$ 7.69
qLogNEI Optimization 4.97 $\pm$ 0.18 4.44 $\pm$ 0.05 4.35 $\pm$ 0.06
Table 1: Total runtime comparison (in seconds) for HIPE, NIPV, and Random across different components over the course of a 4-batch, $q=8$ optimization run. HIPE is more time-consuming to optimize compared to NIPV. While NIPV takes approximately half as long to optimize, the runtime to optimize HIPE is small relative to the time required to fit the fully Bayesian GP with NUTS.
5.5 High-Dimensional SVM Tasks
Finally, we evaluate HIPE and baseline methods on challenging high-dimensional SVM hyperparameter optimization tasks with $D=20$ and $D=40$ input dimensions, considered in similar variants by Ament et al. (2023a); Eriksson and Jankowiak (2021); Papenmeier et al. (2023); Hvarfner et al. (2024). For both tasks, only the last two dimensions—corresponding to the SVM’s global regularization parameters—significantly influence the objective, while the remaining dimensions, corresponding to feature-specific lengthscales, are of lesser importance. Effectively identifying and prioritizing these relevant dimensions is critical for successful optimization. We again consider the two-shot setting, using a larger batch size of $q=32$ due to the higher dimensionality of the problems.
The left panel of Fig. 5 reports the out-of sample inference performance after each batch. On the 2020D task, HIPE achieves competitive, mid-range performance relative to the evaluated methods. On the more challenging 4040D task, HIPE obtains the highest performance, albeit by a narrow margin over the next-best alternatives. The limited budget relative to the dimensionality presents a significant challenge, as the ability to accurately learn the model hyperparameters diminishes with increasing dimensionality. Despite this, we observe in the right panel of Fig. 5 that HIPE identifies the important hyperparameters remarkably well after initialization on the 4040D task—assigning the last two dimensions lengthscales that are, on average, nearly half an order of magnitude smaller than those inferred under a Sobol initialization. In Fig.10, we display the inferred hyperparameter values for NIPV and Random as well. While these methods infer lower lengthscales for the last two dimensions than Sobol, they do not manage to infer as low values as HIPE.
Finally, we note that the best solutions to the SVM problem were almost exclusively located near the boundaries of the search space—particularly in the last two dimensions, which neither NIPV nor HIPE naturally explore during initialization.
Figure 5: Results on $\mathbf{20}$D and $\mathbf{40}$D SVM hyperparameter optimization tasks. Left: Objective value of inferred maximum after each batch. On 2020D, HIPE achieves average performance, climbing to a mid-tier result in the second batch after focusing on hyperparameter learning in the first batch. HIPE obtains a large standard error in the first batch, as two repetitions poorly infers the maximizer and suggests ill-performing points as a result. On 4040D, it outperforms all baselines, demonstrating strong robustness in higher dimensions and the ability to recover after a less successful first batch. Notably, the performance of Random decreases between batches on the 20D task, demonstrating the difficulty of accurately inferring the optimum. Right: Log-mean estimated lengthscale hyperparameters after initialization on the 4040D task. HIPE identifies the last two dimensions–corresponding to the SVM’s global regularization parameters – much more effectively than Sobol, assigning substantially smaller lengthscales to the relevant inputs.
6 Discussion
Contributions
We introduced HIPE, a principled, hyperparameter-free information-theoretic method for initializing Bayesian Optimization and Bayesian Active Learning algorithms. its HIPE yields initial designs that balance coverage of (relevant) areas of the domain with the ability to effectively learn model hyperparameters. HIPE is especially useful in the few-shot, large-batch setting, where it achieves superior surrogate model quality compared to various initialization baselines, as demonstrated by our experiments.
Limitations
HIPE can become computationally expensive, especially for large batch sizes $q$ in higher dimensions $D$, though in many applications this cost is still insignificant compared to the time and resources required to evaluate the underlying black-box function. Finally, the current paper focuses on GP surrogates, and while our main insights and the general approach apply also to other surrogate types, our implementation does not translate directly. Thus, the computation of the acquisition function would have to be re-derived for other types of Bayesian models.
Future work
Here we studied the “cold start” problem of initializing Bayesian Optimization and Bayesian Active Learning from scratch. In practice, we may have access to data from related (but not necessarily identical) problems. This motivates an extension of HIPE to the transfer learning setting, e.g., by means of using a multi-task GP surrogate. Additionally, incorporating prior knowledge of domain experts in a principled fashion is of high practical relevance, and can readily be utilized in the form of a non-uniform $p_{*}$. Finally, we are also interested in studying the multi-objective setting in which different surrogates of potentially different form with different hyperparameters and priors model different objectives but share observations at the same input locations.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: All claims made in the abstract and in the introduction are detailed and backed up with theoretical and empirical results. The content in the paper does not go beyond the broad scope of what is claimed in abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations such as computational scalability are discussed in Section 6 and are acknowledged throughout the paper.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The only theoretical contribution in this paper is Proposition 1, for which assumptions are discussed and a proof is provided in Appendix B.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Implementation details on our main contribution are discussed in the paper, the experimental setup is detailed in the Appendix, and our results can easily be reproduced using the accompanying code submission.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our code submission allows to easily reproduce the results in the paper. Furthermore, our main contribution, HIPE, will be made available in the popular open source library BoTorch (https://github.com/pytorch/botorch) upon publication.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Details on the experimental setup are provided in Appendix A as well as part of the code submission.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All experimental results provide confidence intervals and experiments have been performed with sufficiently many replications to highlight any claimed differences between our contributions and baselines.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The compute resources used for the experiments in the paper are summarized in Appendix A.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research done for this paper neither involved any human subjects nor did it generate or use any data sets that could raise potential privacy concerns. Moreover, since it focuses on foundational methodological contributions to Bayesian Optimization and Bayesian Active Learning, it is highly unlikely to generate any adverse societal impact and potential harmful consequences.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This paper focuses on foundational methodological contributions to Bayesian Optimization and Bayesian Active Learning. As such, the contributions are highly unlikely to generate any adverse societal impact and potential harmful consequences.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This work does not involve the release of any data or models that would pose any risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Work on this paper did not require the use of any models. Any code or data sets that were used are properly credited and their respective licenses are mentioned and properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The code submission includes documentation both in the form of a readme file as well as docstrings throughout the code.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs in any way.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Experimental Setup
We describe the full experimental setup used in the paper: the design of the Bayesian optimization and active learning loops, the benchmarks used, the compute as well as the licenses of all software and datasets. Our code, including all the benchmarks and plotting to reproduce our results, is available at https://github.com/hipeneurips/HIPE.
A.1 Bayesian Optimization Loop
All experiments were conducted using a standardized pipeline based on BoTorch (Balandat et al., 2020) and GPyTorch (Gardner et al., 2018). We use fully Bayesian Gaussian Process (GP) models, with hyperparameter inference performed via the No-U-Turn Sampler (NUTS )(Hoffman and Gelman, 2014) as implemented in Pyro (Bingham et al., 2018). Unless otherwise specified, we draw 192 burn-in samples followed by 288 hyperparameter samples, retaining every 24th sample for evaluation.
Our GP prior is adapted from Hvarfner et al. (2024) to better suit a fully Bayesian setting. Specifically, we set $\mu_{0}=-0.75$ and $\sigma=0.75$, resulting in $\ell_{d}\sim\mathcal{LN}(0.75+\log(D)/2,0.75)$. The noise standard deviation is modeled as $\sigma_{\varepsilon}\sim\mathcal{LN}(-5.5,0.75)$, and the constant mean parameter follows $c\sim\mathcal{N}(0,0.25)$. These modifications ensure that the means of the priors for $\ell_{d}$ and $\sigma_{\varepsilon}^{2}$ approximately match the modes of the corresponding parameters in the prior proposed by Hvarfner et al. (2024), producing similar hyperparameter values in practice.
Acquisition functions are optimized jointly over the batch using multi-start L-BFGS-B optimization with 4 random restarts and 384 initial samples drawn from a scrambled Sobol sequence. For the optimization of LogEI, we additionally sample 384 points from a Gaussian distribution centered at the current incumbent to improve local search performance.
For our proposed HIPE method and relevant baselines (BALD and NIPV), Monte Carlo estimators use $M=12$ hyperparameter samples, $T=1024$ test points drawn uniformly from the search space, and $N=128$ predictive posterior samples.
A.2 Benchmarks
We use three types of benchmarks: synthetic optimization test functions, surrogate-based hyperparameter optimization tasks from LCBench (Zimmer et al., 2021), and high-dimensional SVM hyperparameter optimization problems. Synthetic functions are standard benchmarks for evaluating active learning and Bayesian optimization under controlled noise and dimensionality. LCBench tasks are GP surrogate models trained on 2,000 evaluations of multi-layer perceptrons (MLPs) on real-world datasets, using a Matern 3/2 kernel; all surrogates are included as part of our code release. The SVM benchmarks follow the setup in Ament et al. (2023a), based on the problem originally introduced in Eriksson and Jankowiak (2021). For these tasks, a Support Vector Regressor (Drucker et al., 1996) is fit to the CTSlice dataset, using a fixed subset of 5,000 data points, with 20% reserved for validation. To vary the dimensionality, an XGBoost model is fit to the original 388-dimensional dataset, and the most important features are retained according to XGBoost’s default feature selection criterion. The objective is to minimize the validation RMSE.
Table 2: Noise levels for all Active Learning (AL) and Bayesian Optimization (BO) benchmarks.
Category Benchmark Task Type Dimensionality $\sigma_{\varepsilon}$
Synthetic Ackley (4D) BO 4 2.0
Hartmann (6D) BO / AL 6 0.5
Hartmann (6D) BO / AL 12 0.5
Hartmann (4D) BO 4 0.5
Hartmann (4D) BO 8 0.5
LCBench Car AL 7 2.5
Australian AL 7 2.5
Fashion-MNIST BO 7 0.0
MiniBooNE BO 7 0.0
Car BO 7 0.0
Higgs BO 7 0.0
Segment BO 7 0.0
SVM Feature-reduced SVM BO 20 0.0
Feature-reduced SVM BO 40 0.0
In synthetic functions, additive Gaussian noise is introduced directly to the function evaluations. The LCBench benchmarks rely on pre-trained surrogate models, where the posterior mean is evaluated, and in cases where $\sigma_{\varepsilon}$ is non-zero, additional noise is added to the evaluation of the posterior mean.
A.3 Compute Resources
All experiments were conducted using an NVIDIA A40 GPU cluster. The compute usage to run all experiments in the main paper amounts to approximately 1000 GPU hours, and an additional 500 GPU hours to produce all the results provided in Appendix C.
A.4 Licenses
The following software packages, libraries and datasets were used in our experiments and for presenting the results in the paper:
  • GPyTorch, BoTorch, Hydra: MIT License
  • PyTorch, NumPy, SciPy, Pandas: BSD Licenses
  • Matplotlib, Seaborn: PSF/BSD Licenses
  • LCBench: Apache License
Appendix B Derivation of HIPE as Joint Information Gain
Recall Proposition 1 from section 4.1. See 1
Proof of Proposition 1.
For $\beta=1$, we have
$\textsc{HIPE}_{1}({\bm{X}})$ $=-\mathbb{E}_{\bm{\theta},y({\bm{X}})}\left[\mathbb{E}_{\bm{x}_{*}}\left[\text{H}\left[y(\bm{x}_{*})\mid\bm{\theta},y({\bm{X}})\right]\right]\right]+\left(\text{H}[y({\bm{X}})]-\mathbb{E}_{\bm{\theta}}\left[\text{H}[y({\bm{X}})\mid\bm{\theta}]\right]\right)$
Therefore, with ${\bm{X}}^{*}:=\mathop{\mathrm{arg\,max}}_{{\bm{X}}\in\mathbb{R}^{q\times D}}\textsc{HIPE}_{1}({\bm{X}})$,
${\bm{X}}^{*}$ $=\mathop{\mathrm{arg\,max}}_{{\bm{X}}\in\mathbb{R}^{q\times D}}-\mathbb{E}_{\bm{\theta},y({\bm{X}})}\left[\mathbb{E}_{\bm{x}_{*}}\left[\text{H}\left[y(\bm{x}_{*})\mid\bm{\theta},y({\bm{X}})\right]\right]\right]+\left(\text{H}[y({\bm{X}})]-\mathbb{E}_{\bm{\theta}}\left[\text{H}[y({\bm{X}})\mid\bm{\theta}]\right]\right)$ (11a)
$=\mathop{\mathrm{arg\,max}}_{{\bm{X}}\in\mathbb{R}^{q\times D}}\;-\mathbb{E}_{\bm{\theta},y({{\bm{X}}})}\left[\mathbb{E}_{\bm{x}_{*}}[\text{H}[y(\bm{x}_{*})|\bm{\theta},y({\bm{X}})]]\right]-\mathbb{E}_{\bm{\theta}}[\text{H}[\bm{\theta}|y({\bm{X}})]]$ (11b)
$=\mathop{\mathrm{arg\,max}}_{{\bm{X}}\in\mathbb{R}^{q\times D}}\;\mathbb{E}_{\bm{x}_{*}}\left[\text{H}[y(\bm{x}_{*}),\bm{\theta})]-\mathbb{E}_{y({\bm{X}})}[\text{H}[y(\bm{x}_{*}),\bm{\theta}|y({\bm{X}})]]\right]\;\;\texttt{(by def)}$ (11c)
$=\mathop{\mathrm{arg\,max}}_{{\bm{X}}\in\mathbb{R}^{q\times D}}\;\mathbb{E}_{\bm{x}_{*}}[\textsc{EIG}(y(\bm{x}_{*}),\bm{\theta};{\bm{X}})]$ (11d)
where the equalities follow from the Bayes’ rule of conditional entropy and the fact that the $\mathop{\mathrm{arg\,max}}$ is independent of quantities which do not involve $\bm{X}$}. ∎
Appendix C Additional Experiments
We present supplementary experiments to further analyze the behavior of the evaluated methods under varying experimental conditions. These results provide additional insights into the robustness of the methods with respect to batch size during initialization and active learning, as well as their performance trade-offs across different evaluation metrics.
C.1 Impact of Batch Size on Predictive Performance
We investigate how the initialization batch size $q$ affects predictive quality and inference performance on the LCBench tasks introduced in Section 5.3. We vary $q$ from 6 to 24 in increments of 2 up to $q=16$, then in steps of 4. For each batch size, we evaluate models trained on the initialization batch only — before any active learning iterations — under a fixed-noise setting. This isolates the impact of the initial design and avoids the confounding effects of subsequent data acquisition.
Fig. 6 summarizes the results across three key metrics: test-set NLL, test-set RMSE, and out-of-sample inference performance. HIPE consistently achieves the lowest NLL across most benchmarks and batch sizes, particularly in the $q\in[12,16]$ range, demonstrating strong calibration and hyperparameter learning. In contrast, RMSE results show that HIPE often underperforms relative to simpler baselines like Sobol and Random, suggesting a trade-off between predictive calibration and pointwise accuracy. Inference performance shows HIPE leading for smaller batch sizes ($q<16$), although performance plateaus or declines at higher $q$, indicating that additional samples are not necessarily helpful for inference when uncertainty in relevant parameters remains high. This occurs because, as batch size increases for our experiments, the inferred $\mathop{\mathrm{arg\,max}}$ has an increased propensity to be located at a boundary. As these boundary points may not be well-performing, inference performance occasionally stagnates or decreases with increasing batch size. Notably, as HIPE’s initialization is generally more centered than competing methods’, it obtains higher in-sample values across most batch sizes.
Unlike the active learning setup, these tasks are noiseless and assume fixed, known noise levels during training. This removes uncertainty in the noise model and creates a distinct evaluation setting focused solely on input selection and hyperparameter inference.
Figure 6: Effect of initialization batch size $q$ on predictive quality and inference across LCBench tasks. Each row shows performance on one of three metrics after the initialization batch: (top) NLL, (middle) RMSE, and (bottom) out-of-sample inference. HIPE consistently leads in NLL and small-$q$ inference, while other methods achieve lower RMSE, indicating a trade-off between model calibration and pointwise prediction accuracy.
C.2 Bayesian Optimization with Different Batch Sizes
We further investigate the impact of batch size on Bayesian optimization performance by evaluating all methods on the LCBench tasks with a reduced batch size of $q=8$. As shown in Fig. 6, HIPE maintains its position as the best-performing method overall, achieving the lowest inference regret across the majority of tasks. Notably, HIPE benefits more than competing methods from identifying high-valued points during the initialization phase, which can be attributed to HIPE biasing towards the center of the search space. However, HIPE maintains its advantage throughout, demonstrating its impact on subsequent BO iterations.
Figure 7: Inference regret per method using a batch size $q$ of 8 across LCBench tasks. HIPE is once again the best-performing method overall, but benefits from finding high-valued points during initialization to a greater extent than other methods.
C.3 Active Learning with Different Batch Sizes
We analyze the effect of batch size in the active learning setting by comparing model performance under small batches ($q=8$) and large batches ($q=24$). This evaluation assesses each method’s robustness to changes in batch size and its capability to maintain predictive accuracy and effective hyperparameter learning under varying evaluation budgets.
As shown in Fig. 8, for smaller batches of $q=8$, HIPE achieves the best performance on the NLL metric and significantly outperforms BALD—the second-best method on NLL—in terms of MSE. In this setting, NIPV and BALD exhibit inconsistent behavior, alternating between the worst performance on MLL and RMSE, respectively.
When the batch size increases to $q=24$, as visualized in Fig. 8, HIPE continues to perform strongly, consistently ranking among the top two methods across both evaluation metrics. Notably, its relative performance improves with larger batch sizes, particularly in terms of RMSE, indicating that HIPE scales more effectively with increased parallelism in query selection. This suggests that HIPE is better suited for scenarios requiring efficient learning under larger batch evaluations.
Figure 8: Model accuracy results in the batch active learning setting with smaller batches ($q=8$). RMSE is reported across various synthetic and LCBench surrogate tasks over 8 batches, using 100100 random seeds per benchmark. HIPE achieves the best performance on NLL and outperforms BALD, the second-best method on NLL, in terms of RMSE. NIPV and BALD alternate in exhibiting the worst performance on MLL and RMSE, respectively.
Figure 9: Model accuracy results in the batch active learning setting with larger batches ($q=24$). RMSE is reported across various synthetic and LCBench surrogate tasks over 3 batches, using 100100 random seeds per benchmark. HIPE consistently ranks among the top two methods across both metrics, achieving a strong balance between hyperparameter learning and predictive accuracy. With larger batch sizes, HIPE exhibits improved relative performance, particularly on RMSE, demonstrating effective scalability under increased parallel evaluations.
C.4 Complementary Hyperparameter Visualization on SVM
We display the inferred hyperparameter values of the left-out methods (Random, NIPV) on the 40D-SVM. Both methods infer more accurate hyperparameter values than Sobol, but less accurate than HIPE. Compared to Sobol, both Random and NIPV assign smaller lengthscales to the relevant dimensions, indicating improved identification of important features. However, their estimates remain less precise than those obtained by HIPE, which more effectively distinguishes the relevant dimensions, namely the last two.
Figure 10: Estimated lengthscale hyperparameters for Sobol, Random, NIPV and HIPE on the 40D-SVM task. Random and NIPV infer low lengthscales for the last two dimensions, which are known to be the most important. However, HIPE does so with greater effectiveness.

### v2

@@
 1 Introduction
-Bayesian Optimization (BO) (Mockus et al., 1978; Jones et al., 1998; Frazier, 2018; Garnett, 2023) is a principled framework for sample-efficient global optimization of black-box functions with applications across diverse fields such as biological discovery (Griffiths and Hernández-Lobato, 2020; Stanton et al., 2022), materials science (Frazier and Wang, 2016; Attia et al., 2020; Ament et al., 2023b), online A/B testing (Agarwal et al., 2018; Letham et al., 2019; Feng et al., 2025), and machine learning hyperparameter optimization (HPO) (Snoek et al., 2014; Feurer et al., 2015).
+Bayesian Optimization (BO) (Mockus1978; jonesei; frazier2018tutorial; garnett_bayesoptbook_2023) is a principled framework for sample-efficient global optimization of black-box functions with applications across diverse fields such as biological discovery (griffiths2020constrained; pmlr-v162-stanton22a), materials science (frazier2016bayesian; attia2020closed; ament2023sustainableconcretebayesianoptimization), online A/B testing (agarwal2018online; letham2019constrained; feng2025fastandslow), and machine learning hyperparameter optimization (HPO) (pmlr-v32-snoek14; Feurer2015InitializingBH).
 BO combines a probabilistic surrogate model—commonly a Gaussian Process (GP)—with an acquisition function to select where to evaluate the unknown objective function.
 In many applications, the runtime of a single black-box function evaluation may restrict the experimenter to a small number of batches – the number of sequential rounds of experiments – but many real-world experimental setups permit conducting multiple experiment simultaneously (e.g., on a parallel compute cluster, in a randomized controlled trial, or batch-testing multiple specimens in a lab).
-The success of Bayesian Optimization in practice is highly sensitive to the quality of the surrogate model (Eriksson and Jankowiak, 2021; Hvarfner et al., 2023; Foldager et al., 2023).
+The success of Bayesian Optimization in practice is highly sensitive to the quality of the surrogate model (eriksson2021high; hvarfner2023selfcorrecting; pmlr-v216-foldager23a).
 This is a challenge particularly during the early stages of optimization when few observations are available.
-In these early stages, a small set of inputs is typically selected at random, or via space-filling or quasi-random sampling strategies such as Latin Hypercube Sampling (LHS) or scrambled Sobol’ sequences (Bossek et al., 2020; Owen, 2023).
-While such strategies aim to achieve broad coverage of the input space for initialization of the surrogate model, they may not necessarily result in improving predictive accuracy, or reduce predictive variance, across the entire input space Ren and Sweet (2024).
-Moreover, they neglect a a second crucial aspect of modeling: the need to accurately infer the model’s hyperparameters (Zhang et al., 2019), such as the kernel lengthscales of a GP.
-With accurate hyperparameter estimates, variation in unimportant dimensions will have less influence on the selection of points in subsequent iterations of BO, leading to more sample-efficient optimization (Eriksson and Jankowiak, 2021; Hvarfner et al., 2023; Müller et al., 2023).
-Conversely, poor hyperparameter estimation may cause subsequent BO iterations to fail to make meaningful progress (Berkenkamp et al., 2019), which may be caused by misidentifying signal for noise, exploring irrelevant dimensions, or returning poor terminal recommendations (Hvarfner et al., 2022).
+In these early stages, a small set of inputs is typically selected at random, or via space-filling or quasi-random sampling strategies such as Latin Hypercube Sampling (LHS) or scrambled Sobol’ sequences (10.1145/3377930.3390155; owen2023practicalqmc).
+While such strategies aim to achieve broad coverage of the input space for initialization of the surrogate model, they may not necessarily result in improving predictive accuracy, or reduce predictive variance, across the entire input space ren2024optimalinitializationbatchbayesian.
+Moreover, they neglect a a second crucial aspect of modeling: the need to accurately infer the model’s hyperparameters (zhang2021ddd), such as the kernel lengthscales of a GP.
+With accurate hyperparameter estimates, variation in unimportant dimensions will have less influence on the selection of points in subsequent iterations of BO, leading to more sample-efficient optimization (eriksson2021high; hvarfner2023selfcorrecting; pmlr-v202-muller23a).
+Conversely, poor hyperparameter estimation may cause subsequent BO iterations to fail to make meaningful progress (berkenkamp2019noregret), which may be caused by misidentifying signal for noise, exploring irrelevant dimensions, or returning poor terminal recommendations (hvarfner2022joint).
 In this work, we provide a principled approach to addressing this initialization challenge. Our main contributions are as follows:
@@
 In practice, the mean function is often kept constant, leaving the covariance function to capture the structural properties of the objective.
-To model differences in variable importance in GPs with stationary kernels, each input dimension is commonly scaled by a lengthscale hyperparameter $\ell_{i}$, a practice known as Automatic Relevance Determination (ARD) (Williams and Rasmussen, 1995).
+To model differences in variable importance in GPs with stationary kernels, each input dimension is commonly scaled by a lengthscale hyperparameter $\ell_{i}$, a practice known as Automatic Relevance Determination (ARD) (NIPS1995_7cce53cf).
 Additional, optional hyperparameters include a learnable noise variance $\sigma_{\varepsilon}^{2}$ and signal variance $\sigma_{f}^{2}$.
 The full set of hyperparameters $\bm{\theta}=\{\bm{\ell},\sigma_{\varepsilon}^{2},\sigma_{f}^{2}\}$ may be learned either by maximizing the marginal likelihood $p(\mathcal{D}\mid\bm{\theta})$ (Maximum Likelihood Estimation, MLE) or by incorporating hyperpriors $p(\bm{\theta})$ to perform Maximum A Posteriori (MAP) estimation.
-Alternatively, a fully Bayesian treatment (Osborne, 2010; Lalchand and Rasmussen, 2020) integrates over $\bm{\theta}$ to approximate the full Bayesian posterior distribution using Markov Chain Monte Carlo (MCMC) methods, thereby explicitly accounting for hyperparameter uncertainty.
-For additional background on GPs, see (Rasmussen and Williams, 2005).
+Alternatively, a fully Bayesian treatment (osborne2010bayesian; lalchand2020approximate) integrates over $\bm{\theta}$ to approximate the full Bayesian posterior distribution using Markov Chain Monte Carlo (MCMC) methods, thereby explicitly accounting for hyperparameter uncertainty.
+For additional background on GPs, see (10.5555/1162254).
 2.2 Bayesian Optimization
-Bayesian Optimization (BO) (Frazier, 2018) is a sample-efficient framework for finding the maximizer $\bm{x}_{*}=\mathop{\mathrm{arg\,max}}_{\bm{x}\in\mathcal{X}}f(\bm{x})$ of a black-box function $f:\mathcal{X}\rightarrow\mathbb{R}$ over a $D$-dimensional input space $\mathcal{X}=[0,1]^{D}$.
+Bayesian Optimization (BO) (frazier2018tutorial) is a sample-efficient framework for finding the maximizer $\bm{x}_{*}=\mathop{\mathrm{arg\,max}}_{\bm{x}\in\mathcal{X}}f(\bm{x})$ of a black-box function $f:\mathcal{X}\rightarrow\mathbb{R}$ over a $D$-dimensional input space $\mathcal{X}=[0,1]^{D}$.
 The function $f$ is assumed to be expensive to evaluate and observable only through noisy point-wise measurements, $y(\bm{x})=f(\bm{x})+\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2})$.
@@
 Acquisition functions balance exploration and exploitation typically through greedy heuristics.
-Popular examples include Expected Improvement (EI) (Jones et al., 1998; Bull, 2011) and its numerically stable variant, LogEI (Ament et al., 2023a), as well as the Upper Confidence Bound (UCB) (Srinivas et al., 2012, 2010).
+Popular examples include Expected Improvement (EI) (jonesei; JMLR:v12:bull11a) and its numerically stable variant, LogEI (ament2023unexpected), as well as the Upper Confidence Bound (UCB) (Srinivas_2012; srinivas2010ucb).
 In batch BO, multiple points are selected in parallel to accelerate data collection.
-This is often achieved by computing and jointly optimizing a (quasi-)MC estimate of the utility $u$ associated with acquisition function over the full batch ${\bm{X}}=\{\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{q}\}$ of size $q$ (Wilson et al., 2017, 2020; Balandat et al., 2020).
+This is often achieved by computing and jointly optimizing a (quasi-)MC estimate of the utility $u$ associated with acquisition function over the full batch ${\bm{X}}=\{\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{q}\}$ of size $q$ (wilson2017reparam; wilson2020efficiently; balandat2020botorch).
 2.3 Bayesian Active Learning
-Bayesian Active Learning (BAL) and Bayesian Experimental Design (BED) (Chaloner and Verdinelli, 1995) aims to improve predictive models by strategically selecting data points that are most informative, either with regard to the model or to future predictions.
-A central quantity is the Expected Information Gain (EIG):
+Bayesian Active Learning (BAL) and Bayesian Experimental Design (BED) (chaloner1995bed) aim to improve predictive models by selecting data points that are most informative, either with regard to the model or to future predictions.
+A key quantity is the Expected Information Gain (EIG):
 $\textsc{EIG}(\xi;y(\bm{x}))=\text{H}[\xi]-\mathbb{E}_{y(\bm{x})}\left[\text{H}[\xi|y(\bm{x})]\right],$ (1)
@@
 Bayesian Active Learning by Disagreement
-Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011; Kirsch et al., 2019) selects query points that maximize the mutual information between model predictions and hyperparameters $\bm{\theta}$:
+Bayesian Active Learning by Disagreement (BALD) (houlsby2011bald; kirsch2019batchbald) selects query points that maximize the mutual information between model predictions and hyperparameters $\bm{\theta}$:
 $\text{BALD}(\bm{x})=\textsc{EIG}(y(\bm{x});\bm{\theta})=\text{H}[y(\bm{x})|\mathcal{D}]-\mathbb{E}_{\bm{\theta}}[\text{H}[y(\bm{x})|\mathcal{D},\bm{\theta}]].$ (2)
@@
 In the GP setting, this often leads to axis-aligned queries when there is high uncertainty in the lengthscales, and to repeated queries when observation noise is highly uncertain.
-Notably, BALD is model-agnostic and can be applied to a wide range of surrogate models and hyperparameters, including subspace models (Garnett et al., 2014) and additive decompositions (Hvarfner et al., 2023; Gardner et al., 2017).
-Negative Integrated Posterior Variance and Expected Predictive Information Gain
-The Negative Integrated Posterior Variance (NIPV) (Chen and Zhou, 2014; Seo et al., 2000) criterion selects queries that minimize the expected posterior variance over a test distribution $p_{*}(\bm{x})$:
+Notably, BALD is model-agnostic and can be applied to a wide range of surrogate models and hyperparameters, including subspace models (garnett2014active) and additive decompositions (hvarfner2023selfcorrecting; pmlr-v54-gardner17a).
+Negative Integrated Posterior Variance
+The Negative Integrated Posterior Variance (NIPV) (sambu2000GPRactiveselection) criterion selects queries that minimize the expected posterior variance over a test distribution $p_{*}(\bm{x})$:
 $\text{NIPV}(\bm{x};p_{*})=-\mathbb{E}_{\bm{x}_{*}\sim p_{*}}\left[\sigma^{2}(\bm{x}_{*})\;|\ \bm{x},\mathcal{D}\right].$ (3)
-Similarly, Expected Predictive Information Gain (EPIG) (Bickford Smith et al., 2023) selects queries that minimize the expected predictive entropy over $p_{*}(\bm{x})$:
-$\text{EPIG}(\bm{x};p_{*})=-\mathbb{E}_{\bm{x}_{*}\sim p_{*}}\left[\text{H}[y(\bm{x})]\;|\;\bm{x}_{*},\mathcal{D}\right].$ (4)
-Both objectives promote the selection of data that reduces uncertainty over the test distribution, but without considering the effect on hyperparameter learning.
+Expected Predictive Information Gain
+The Expected Predictive Information Gain (EPIG) (pmlr-v206-bickfordsmith23a) selects a candidate point $\bm{x}$ by maximizing the mutual information between its label $y(\bm{x})$ and the label $y(\bm{x}^{*})$ at a randomly drawn test input $\bm{x}_{*}\sim p_{*}(\bm{x})$,
+$\mathrm{EPIG}(\bm{x};p_{*})=\mathbb{E}_{\bm{x}_{*}\sim p_{*}}\!\left[I\!\big(y(\bm{x});y(\bm{x}_{*})\mid\mathcal{D}\big)\right].$ (4)
+which can equivalently be expressed in terms of an expectation over a difference of entropies.
+In its general, model-agnostic form, this constitutes of a nested expectation over both test inputs $\bm{x}_{*}$ and outcomes and selected locations $y(\bm{x})$:
+$\mathrm{EPIG}(\bm{x};p_{*})=\mathbb{E}_{\bm{x}_{*}\sim p_{*}}\!\left[\text{H}\!\left[y(\bm{x}_{*})\mid\mathcal{D}\right]-\mathbb{E}_{y(\bm{x})\mid\mathcal{D}}\!\left[\text{H}\!\left[y(\bm{x}_{*})\mid\mathcal{D},\,\bm{x},\,y(\bm{x})\right]\right]\right].$ (5)
+EPIG for Gaussian Processes
+Standard GP regression with Gaussian likelihoods constitutes a special case for EPIG, as the predictive variance (and therefore the predictive entropy) after conditioning on $(\bm{x},y(\bm{x}))$ is independent of the realized value $y(\bm{x})$.
+Thus,
+$\text{H}\!\left[y(\bm{x}_{*})\mid\mathcal{D},\,\bm{x},\,y(\bm{x})\right]=\text{H}\!\left[y(\bm{x}_{*})\mid\mathcal{D},\,\bm{x}\right],$
+and the inner expectation over $y(\bm{x})$ in (5) disappears. EPIG therefore simplifies to an expected reduction in predictive entropy:
+$\mathrm{EPIG}_{\mathrm{GP}}(\bm{x};p_{*})=\mathbb{E}_{\bm{x}_{*}\sim p_{*}}\!\left[\text{H}\!\left[y(\bm{x}_{*})\mid\mathcal{D}\right]-\text{H}\!\left[y(\bm{x}_{*})\mid\mathcal{D},\,\bm{x}\right]\right].$ (6)
+We will use this GP-specific form of EPIG throughout the remainder of this paper.
+Both NIPV and EPIG promote the selection of data that reduces uncertainty over the test distribution, but without considering the effect on hyperparameter learning.
 The test distribution encodes how much emphasis is put on different parts of the domain.
@@
 Initialization has received surprisingly little attention in the context of BO.
-Bossek et al.
-(2020) conducts a study on the effect of various random initial designs on BO performance.
-Alternatively, minimax or maximin criteria (Johnson et al., 1990) may be used to accomplish evenly distributed designs.
-Maybe closest to our work is (Zhang et al., 2019), which proposes LHS-Beta, an initial design criterion which alters samples drawn by LHS to achieve pairwise distances between points which matches a Beta distribution.
+10.1145/3377930.3390155 conducts a study on the effect of various random initial designs on BO performance.
+Alternatively, minimax or maximin criteria (JOHNSON1990131) may be used to accomplish evenly distributed designs.
+Maybe closest to our work is (zhang2021ddd), which proposes LHS-Beta, an initial design criterion which alters samples drawn by LHS to achieve pairwise distances between points which matches a Beta distribution.
 LHS-Beta pursues diverse pairwise distances in the data, in order to best learn the lengthscale of a GP with an isotropic kernel.
-Zimmerman (2006); Müller and Zimmerman (1999) address the problem of learning parameters of Kriging estimators using a empirical estimates of optimal experimental design criteria, limiting candidates to a fixed grid of points.
-Entropy-maximizing (Guestrin et al., 2005; MacKay, 1992, 1995) or variance-minimizing (Park et al., 2024) designs have been explored in active learning for optimal sensor placement (Krause et al., 2006, 2008) and other applications involving GPs, such as geostatistics (Sauer et al., 2023) and contour finding (Cole et al., 2023).
-Moreover, parameter-related EIG criteria are a bedrock of the broader topic of BED (Bickford Smith et al., 2023; Rainforth et al., 2024; Kirsch et al., 2021; Chaloner and Verdinelli, 1995), which focuses on selecting data that is most informative about model parameters or future predictions.
+zimmerman2006optimal; muller1999optimal address the problem of learning parameters of Kriging estimators using a empirical estimates of optimal experimental design criteria, limiting candidates to a fixed grid of points.
+Entropy-maximizing (guestrin2005near; mackay1992information; MACKAY199573) or variance-minimizing (park2024activelearningpiecewisegaussian) designs have been explored in active learning for optimal sensor placement (krause2006near; JMLR:v9:krause08a) and other applications involving GPs, such as geostatistics (Sauer02012023) and contour finding (Cole01012023).
+Moreover, parameter-related EIG criteria are a bedrock of the broader topic of BED (pmlr-v206-bickfordsmith23a; rainforth2024modern; kirsch2021jepig; chaloner1995bed), which focuses on selecting data that is most informative about model parameters or future predictions.
 These prediction or model-oriented criteria have yet to see widespread use in BO, particularly for initialization.
-However, information-theoretic acquisition functions (Hennig and Schuler, 2012; Hernández-Lobato et al., 2014; Wang and Jegelka, 2017; Moss et al., 2021; Hvarfner et al., 2022; Tu et al., 2022; Neiswanger et al., 2021, 2024) address the optimization problem from an information theoretic perspective, albeit not with a primary focus on initialization or model predictive performance.
-The problem of actively learning model hyperparameters during BO (post initialization) has previously been investigated by Hvarfner et al.
-(2023), who propose a combined BO-BAL framework to actively learn the hyperparameters of the GP along with the optimum, and demonstrate that better-calibrated surrogates significantly enhance BO performance.
-Houlsby et al.
-(2011) proposes BALD, an active learning acquisition functions for hyperparameters in preference learning in GPs.
-Riis et al.
-(2022) proposes a a Query-By-Committee-oriented acquisition function for BAL in GPs.
-Lastly, Berkenkamp et al.
-(2019); Ziomek et al.
-(2024) address BO performance under hyperparameter uncertainty from a theoretical perspective, proving regret bounds when hyperparameters of the objective are unknown.
+However, information-theoretic acquisition functions (entropysearch; pes; wang2017maxvalue; moss2021gibbon; hvarfner2022joint; tu2022joint; pmlr-v139-neiswanger21a; 10.5555/3600270.3601798) address the optimization problem from an information theoretic perspective, albeit not with a primary focus on initialization or model predictive performance.
+The problem of actively learning model hyperparameters during BO (post initialization) has previously been investigated by hvarfner2023selfcorrecting, who propose a combined BO-BAL framework to actively learn the hyperparameters of the GP along with the optimum, and demonstrate that better-calibrated surrogates significantly enhance BO performance.
+houlsby2011bald proposes BALD, an active learning acquisition functions for hyperparameters in preference learning in GPs.
+riis2022bayesian proposes a a Query-By-Committee-oriented acquisition function for BAL in GPs.
+Lastly, berkenkamp2019noregret; NEURIPS2024_9cf904c8 address BO performance under hyperparameter uncertainty from a theoretical perspective, proving regret bounds when hyperparameters of the objective are unknown.
 4 Method
@@
 If the initial design fails to reduce uncertainty in key hyperparameters or leaves large regions of the input space unexplored, both hyperparameter inference and optimization become challenging.
-This setting is common if evaluations are lengthy but can be effectively parallelized, which is the case for instance in online A/B tests or lab experiments for biology Stanton et al.
-(2022); González-Duque et al.
-(2024) or material design Yik et al.
-(2025).
+This setting is common if evaluations are lengthy but can be effectively parallelized, which is the case for instance in online A/B tests or lab experiments for biology pmlr-v162-stanton22a; gonzalez2024survey or material design yik2025battery.
 […]
@@
 In fact, space-filling is not synonymous with informativeness: a model trained on a space-filling design may still exhibit high predictive uncertainty in regions most relevant for optimization.
-For example, random designs increasingly sample near the boundaries of the search space as dimensionality grows Swersky (2017); Köppen (2000), whereas a point located at the center of the search space may be far more informative for optimization, as it can yield a greater reduction in predictive variance throughout the search space.
-To this point, a strategy that explicitly reduces predictive uncertainty within the region of interest Bickford Smith et al.
-(2023)—namely, the search space—not only seems better suited for our goals, but is also better aligned with both theoretical (Srinivas et al., 2010; Berkenkamp et al., 2019) and information gain-based criteria, as well as the practical objective of achieving a more informed, in an informal sense, model post-initialization Garnett (2023).
-Methods like NIPV (Zhang et al., 2019) which minimize predictive variance over the input space achieve this goal better than Sobol, and yields designs that are spread out but not necessarily informative for model calibration and hyperparameter learning.
-On the other hand, methods that focus exclusively on hyperparameter informativeness, such as BALD (Houlsby et al., 2011), tend to cluster queries along specific axes in order to resolve lengthscales, thereby sacrificing coverage and potentially overlooking important regions of the search space.
+For example, random designs increasingly sample near the boundaries of the search space as dimensionality grows swersky2017improving; koppencurse, whereas a point located at the center of the search space may be far more informative for optimization, as it can yield a greater reduction in predictive variance throughout the search space.
+To this point, a strategy that explicitly reduces predictive uncertainty within the region of interest pmlr-v206-bickfordsmith23a—namely, the search space—not only seems better suited for our goals, but is also better aligned with both theoretical (srinivas2010ucb; berkenkamp2019noregret) and information gain-based criteria, as well as the practical objective of achieving a more informed, in an informal sense, model post-initialization garnett_bayesoptbook_2023.
+Methods like NIPV (zhang2021ddd) which minimize predictive variance over the input space achieve this goal better than Sobol, and yields designs that are spread out but not necessarily informative for model calibration and hyperparameter learning.
+On the other hand, methods that focus exclusively on hyperparameter informativeness, such as BALD (houlsby2011bald), tend to cluster queries along specific axes in order to resolve lengthscales, thereby sacrificing coverage and potentially overlooking important regions of the search space.
 When additional hyperparameters are present, such as a learnable noise variance, BALD may even select duplicate queries at the same location to better estimate said noise variance.
@@
 We approach the problem of initialization through the lens of optimization, by simultaneously maximizing the coverage of a region of interest and the information acquired about model-level uncertainty.
-A natural way to achieve this is to optimize a criterion that combines the EPIG and BALD objectives:
+A natural way to achieve this is to optimize a criterion that combines EPIG and BALD:
 $\text{{HIPE}}_{\beta}({\bm{X}}):=\underbrace{-\mathbb{E}_{\bm{\theta},y({\bm{X}})}\left[\mathbb{E}_{\bm{x}_{*}}\left[\text{H}\left[y(\bm{x}_{*})\mid\bm{\theta},y({\bm{X}})\right]\right]\right]}_{\text{EPIG objective}}+\beta\;\underbrace{\left(\text{H}[y({\bm{X}})]-\mathbb{E}_{\bm{\theta}}\left[\text{H}[y({\bm{X}})\mid\bm{\theta}]\right]\right)}_{\text{BALD objective}}$ (7)
@@
 To what extent the reduction in hyperparameter entropy manifests in a reduction in test set entropy can be quantified through the mutual information between the hyperparameters $\bm{\theta}$ and the test points $y(\bm{x}_{*}),\;\bm{x}_{*}\sim p_{*}(\bm{x})$:
-$\textsc{EIG}(y(\bm{x}_{*});\bm{\theta}|\mathcal{D})=\mathbb{E}_{\bm{x}_{*}}\left[\text{H}[y(\bm{x}_{*})|\mathcal{D}]-\mathbb{E}_{\bm{\theta}}[\text{H}[y(\bm{x}_{*})|\bm{\theta},\mathcal{D}]]\right].$ (7)
+$\beta=\textsc{EIG}(y(\bm{x}_{*});\bm{\theta}|\mathcal{D})=\mathbb{E}_{\bm{x}_{*}}\left[\text{H}[y(\bm{x}_{*})|\mathcal{D}]-\mathbb{E}_{\bm{\theta}}[\text{H}[y(\bm{x}_{*})|\bm{\theta},\mathcal{D}]]\right].$ (9)
 Intuitively, Eq.
@@
 4.2 A Parallel Monte Carlo Implementation of HIPE
-Following the Monte Carlo approach used in modern acquisition functions (Wilson et al., 2020; Balandat et al., 2020), we implement a parallel version of HIPE that enables joint optimization.
+Following the Monte Carlo approach used in modern acquisition functions (wilson2020efficiently; balandat2020botorch), we implement a parallel version of HIPE that enables joint optimization.
 For a candidate batch ${\bm{X}}=\{\bm{x}_{1},\dots,\bm{x}_{q}\}$, we estimate the acquisition objective using $M$ MC samples over the hyperparameters, $T$ test locations of the EPIG objective, and $N$ samples from the predictive posterior for the BALD objective:
@@
 Thus, the BALD objective amounts to repeated evaluation the $q$-dimensional multivariate normal posterior $y({\bm{X}})$ for each candidate $\bm{X}$}, and estimating the posterior entropy from it.
-Secondly, we estimate the EPIG objective analogously to Balandat et al.
-(2020) as
+Secondly, we estimate the EPIG objective analogously to balandat2020botorch as
 $\alpha_{\text{EPIG}}({\bm{X}};p_{*},\bm{\theta})\approx\frac{1}{M}\frac{1}{T}\sum_{m=1}^{M}\sum_{t=1}^{T}\left[C-\text{H}[y(\bm{x}_{*}^{(t)})|{\bm{X}},\bm{\theta}^{(m)},\mathcal{D}]\right],$ (12)
@@
 With these two MC estimators, the subsequent optimization can be carried out jointly for the entire $q$-batch in a $qD$-dimensional space, as the entropy reduction is computed with regard to the entire batch of candidates as opposed to a singular one.
-Using Sample Average Approximation (Balandat et al., 2020), the HIPE objective is deterministic and auto-differentiable.
-One downside of this formulation is that the nested MC estimator imposes substantial computational runtime, making HIPE less suited for high-throughput applications (Eriksson et al., 2019; Maus et al., 2022; Daulton et al., 2022).
+Using Sample Average Approximation (balandat2020botorch), the HIPE objective is deterministic and auto-differentiable.
+One downside of this formulation is that the nested MC estimator imposes substantial computational runtime, making HIPE less suited for high-throughput applications (eriksson2019turbo; maus2022lolbo; pmlr-v162-daulton22a).
 However, in those applications the quality of the initialization batch is generally much less crucial, so this is not a limitation in practice.
@@
 We also study how these model quality improvements translate to better optimization performance.
-To this end, we consider the out-of-sample inference performance (Hernández-Lobato et al., 2014; Hvarfner et al., 2022), that is, the performance of the point $\bm{x}^{\prime}=\mathop{\mathrm{arg\,max}}\mu(\bm{x}\mid\mathcal{D})$ selected as the maximizer of the posterior mean of the surrogate model fit on the data available in each batch.
+To this end, we consider the out-of-sample inference performance (pes; hvarfner2022joint), that is, the performance of the point $\bm{x}^{\prime}=\mathop{\mathrm{arg\,max}}\mu(\bm{x}\mid\mathcal{D})$ selected as the maximizer of the posterior mean of the surrogate model fit on the data available in each batch.
 We choose this metric since using observed points directly performs very poorly in noisy settings, and only considering in-sample points is rather limiting in the few-shot setting.
 5.1 Batch Active Learning on GP Surrogates and Synthetic Functions
-We first evaluate the ability of HIPE to learn accurate surrogate models through batch active learning on noisy synthetic test functions and surrogate LCBench (Zimmer et al., 2021) tasks.
-The LCBench tasks are derived from complete neural network training runs on various OpenML (Vanschoren et al., 2014) datasets, with 7D GP surrogate models fitted as described in Appendix A.
-Additionally, we evaluate performance on the Hartmann 66D function and a high-dimensional Hartmann 66D (1212D) variant, where dummy input dimensions are added following standard practice in high-dimensional BO (Eriksson and Jankowiak, 2021).
+We first evaluate the ability of HIPE to learn accurate surrogate models through batch active learning on noisy synthetic test functions and surrogate LCBench (ZimLin2021a) tasks.
+The LCBench tasks are derived from complete neural network training runs on various OpenML (vanschoren2014openml) datasets, with 7D GP surrogate models fitted as described in Appendix A.
+Additionally, we evaluate performance on the Hartmann 66D function and a high-dimensional Hartmann 66D (1212D) variant, where dummy input dimensions are added following standard practice in high-dimensional BO (eriksson2021high).
 These dummy dimensions introduce an additional challenge, as effectively identifying and ignoring irrelevant features is critical for accurate predictions.
@@
 On NLL, HIPE performs comparably to BALD, which targets hyperparameter learning and thus excels at model calibration.
-Similarly, HIPE is competitive with NIPV on RMSE, a metric for which NIPV is particularly well-suited (Gramacy and Lee, 2009).
+Similarly, HIPE is competitive with NIPV on RMSE, a metric for which NIPV is particularly well-suited (gramacy2009adaptive).
 Figure 2: Model accuracy results in the batch active learning setting.
-We report RMSE across various synthetic and LCBench surrogate tasks over 4 batches of $q=16$ evaluations across 100100 seeds per benchmark.
+We report RMSE across various synthetic and LCBench surrogate tasks over 4 batches of $q=16$ evaluations across 100100 seeds.
 HIPE consistently ranks in the top two in relative rankings on both metrics, achieving a strong balance between hyperparameter learning and predictive accuracy.
-BALD performs competitively on marginal log-likelihood (MLL) but underperforms on RMSE due to limited space-filling behavior, while NIPV excels at reducing RMSE but struggles with model calibration.
+BALD performs competitively on marginal log-likelihood (MLL) but underperforms on RMSE due to poor space-filling, while NIPV excels at reducing RMSE but struggles with calibration.
 On aggregate, random initialization methods lag significantly behind across all benchmarks.
@@
 Random and space-filling initialization methods (LHS-Beta, Random, and Sobol) perform noticeably worse across all settings.
+In App.
+C.1, we show the runtime of HIPE and various other methods on Hartmann (6D).
+The extra cost of HIPE acquisition is a minority of iteration runtime for a fully Bayesian model.
 Figure 3: Out-of-sample inference optimization performance on noisy synthetic benchmark functions under the two-shot setting ($B=2$, $q=24$) across 100100 seeds per benchmark.
@@
 7, we demonstrate the performance of the same initialization schemes on $q=8$, $B=5$.
-In this setting, HIPE is the top-performing method on all benchmarks, and the overall ranking of algorithms remains roughly intact.
+In this setting, HIPE remains the top-performing method, and the overall ranking of algorithms remains intact.
 Figure 4: Two-shot optimization results on five hyperparameter optimization tasks from LCBench across 200 seeds for each task.
@@
 This confirms the practical relevance of our approach for real-world HPO scenarios, where both accurate predictive modeling and effective exploration are crucial.
-5.4 Runtime Analysis
-We compare the time required to optimize HIPE to other options, namely NIPV, as well as its runtime compared to other components of a BO loop that employs fully Bayesian GP modeling.
-In Table 1, we demonstrate that, while HIPE is substantially slower to optimize than NIPV, its total runtime is still a minor component of the overall runtime of the BO run.
-Runtime Component (sec) HIPE NIPV Random
-Model fit (initial) 21.91 $\pm$ 0.15 21.91 $\pm$ 0.15 N/A
-Acquisition Opt. 19.61 $\pm$ 1.12 8.62 $\pm$ 0.37
-Model fit (BO loop) 181.79 $\pm$ 7.60 189.90 $\pm$ 5.17 177.33 $\pm$ 7.69
-qLogNEI Optimization 4.97 $\pm$ 0.18 4.44 $\pm$ 0.05 4.35 $\pm$ 0.06
-Table 1: Total runtime comparison (in seconds) for HIPE, NIPV, and Random across different components over the course of a 4-batch, $q=8$ optimization run.
-HIPE is more time-consuming to optimize compared to NIPV.
-While NIPV takes approximately half as long to optimize, the runtime to optimize HIPE is small relative to the time required to fit the fully Bayesian GP with NUTS.
 5.4 High-Dimensional SVM Tasks
-Finally, we evaluate HIPE and baseline methods on challenging high-dimensional SVM hyperparameter optimization tasks with $D=20$ and $D=40$ input dimensions, considered in similar variants by Ament et al.
-(2023a); Eriksson and Jankowiak (2021); Papenmeier et al.
-(2023); Hvarfner et al.
-(2024).
+Finally, we evaluate HIPE and baseline methods on challenging high-dimensional SVM hyperparameter optimization tasks with $D=20$ and $D=40$ input dimensions, considered in similar variants by ament2023unexpected; eriksson2021high; papenmeier2023bounce; hvarfner2024vanilla.
 For both tasks, only the last two dimensions—corresponding to the SVM’s global regularization parameters—significantly influence the objective, while the remaining dimensions, corresponding to feature-specific lengthscales, are of lesser importance.
@@
 its HIPE yields initial designs that balance coverage of (relevant) areas of the domain with the ability to effectively learn model hyperparameters.
-HIPE is especially useful in the few-shot, large-batch setting, where it achieves superior surrogate model quality compared to various initialization baselines, as demonstrated by our experiments.
+HIPE is especially useful in the few-shot, large-batch setting, where it achieves superior surrogate model quality compared to various initialization baselines.
 Limitations
@@
 Finally, we are also interested in studying the multi-objective setting in which different surrogates of potentially different form with different hyperparameters and priors model different objectives but share observations at the same input locations.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4.
-Justification: All claims made in the abstract and in the introduction are detailed and backed up with theoretical and empirical results.
-The content in the paper does not go beyond the broad scope of what is claimed in abstract and introduction.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: Limitations such as computational scalability are discussed in Section 6 and are acknowledged throughout the paper.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: The only theoretical contribution in this paper is Proposition 1, for which assumptions are discussed and a proof is provided in Appendix B.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19.
-Justification: Implementation details on our main contribution are discussed in the paper, the experimental setup is detailed in the Appendix, and our results can easily be reproduced using the accompanying code submission.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24.
-Justification: Our code submission allows to easily reproduce the results in the paper.
-Furthermore, our main contribution, HIPE, will be made available in the popular open source library BoTorch (https://github.com/pytorch/botorch) upon publication.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: Details on the experimental setup are provided in Appendix A as well as part of the code submission.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34.
-Justification: All experimental results provide confidence intervals and experiments have been performed with sufficiently many replications to highlight any claimed differences between our contributions and baselines.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: The compute resources used for the experiments in the paper are summarized in Appendix A.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44.
-Justification: The research done for this paper neither involved any human subjects nor did it generate or use any data sets that could raise potential privacy concerns.
-Moreover, since it focuses on foundational methodological contributions to Bayesian Optimization and Bayesian Active Learning, it is highly unlikely to generate any adverse societal impact and potential harmful consequences.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49.
-Justification: This paper focuses on foundational methodological contributions to Bayesian Optimization and Bayesian Active Learning.
-As such, the contributions are highly unlikely to generate any adverse societal impact and potential harmful consequences.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: This work does not involve the release of any data or models that would pose any risk for misuse.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59.
-Justification: Work on this paper did not require the use of any models.
-Any code or data sets that were used are properly credited and their respective licenses are mentioned and properly respected.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: The code submission includes documentation both in the form of a readme file as well as docstrings throughout the code.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: The core method development in this research does not involve LLMs in any way.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
 Appendix A Experimental Setup
 We describe the full experimental setup used in the paper: the design of the Bayesian optimization and active learning loops, the benchmarks used, the compute as well as the licenses of all software and datasets.
-Our code, including all the benchmarks and plotting to reproduce our results, is available at https://github.com/hipeneurips/HIPE.
+The code to run the experiments in the paper, including all the benchmarks and plotting to reproduce our results, is available at https://github.com/hipeneurips/HIPE.
+An easy-to-run, tested version of the algorithm and accompanying notebook is available open-source in BoTorch.
+The HIPE acquisition is found at https://github.com/meta-pytorch/botorch/blob/main/botorch_community/acquisition/bayesian_active_learning.py.
+The notebook, which demonstrates how to run the algorithm should be run in its indented setting and the expected results, can be found at https://github.com/meta-pytorch/botorch/blob/main/notebooks_community/hyperparameter_informed_predictive_exploration.
 A.1 Bayesian Optimization Loop
-All experiments were conducted using a standardized pipeline based on BoTorch (Balandat et al., 2020) and GPyTorch (Gardner et al., 2018).
-We use fully Bayesian Gaussian Process (GP) models, with hyperparameter inference performed via the No-U-Turn Sampler (NUTS )(Hoffman and Gelman, 2014) as implemented in Pyro (Bingham et al., 2018).
+All experiments were conducted using a standardized pipeline based on BoTorch (balandat2020botorch) and GPyTorch (gardner2018gpytorch).
+We use fully Bayesian Gaussian Process (GP) models, with hyperparameter inference performed via the No-U-Turn Sampler (NUTS )(JMLR:v15:hoffman14a) as implemented in Pyro (bingham2018pyro).
 Unless otherwise specified, we draw 192 burn-in samples followed by 288 hyperparameter samples, retaining every 24th sample for evaluation.
-Our GP prior is adapted from Hvarfner et al.
-(2024) to better suit a fully Bayesian setting.
+Our GP prior is adapted from hvarfner2024vanilla to better suit a fully Bayesian setting.
 Specifically, we set $\mu_{0}=-0.75$ and $\sigma=0.75$, resulting in $\ell_{d}\sim\mathcal{LN}(0.75+\log(D)/2,0.75)$.
 The noise standard deviation is modeled as $\sigma_{\varepsilon}\sim\mathcal{LN}(-5.5,0.75)$, and the constant mean parameter follows $c\sim\mathcal{N}(0,0.25)$.
-These modifications ensure that the means of the priors for $\ell_{d}$ and $\sigma_{\varepsilon}^{2}$ approximately match the modes of the corresponding parameters in the prior proposed by Hvarfner et al.
-(2024), producing similar hyperparameter values in practice.
+These modifications ensure that the means of the priors for $\ell_{d}$ and $\sigma_{\varepsilon}^{2}$ approximately match the modes of the corresponding parameters in the prior proposed by hvarfner2024vanilla, producing similar hyperparameter values in practice.
 Acquisition functions are optimized jointly over the batch using multi-start L-BFGS-B optimization with 4 random restarts and 384 initial samples drawn from a scrambled Sobol sequence.
@@
 A.2 Benchmarks
-We use three types of benchmarks: synthetic optimization test functions, surrogate-based hyperparameter optimization tasks from LCBench (Zimmer et al., 2021), and high-dimensional SVM hyperparameter optimization problems.
+We use three types of benchmarks: synthetic optimization test functions, surrogate-based hyperparameter optimization tasks from LCBench (ZimLin2021a), and high-dimensional SVM hyperparameter optimization problems.
 Synthetic functions are standard benchmarks for evaluating active learning and Bayesian optimization under controlled noise and dimensionality.
 LCBench tasks are GP surrogate models trained on 2,000 evaluations of multi-layer perceptrons (MLPs) on real-world datasets, using a Matern 3/2 kernel; all surrogates are included as part of our code release.
-The SVM benchmarks follow the setup in Ament et al.
-(2023a), based on the problem originally introduced in Eriksson and Jankowiak (2021).
-For these tasks, a Support Vector Regressor (Drucker et al., 1996) is fit to the CTSlice dataset, using a fixed subset of 5,000 data points, with 20% reserved for validation.
+The SVM benchmarks follow the setup in ament2023unexpected, based on the problem originally introduced in eriksson2021high.
+For these tasks, a Support Vector Regressor (svr) is fit to the CTSlice dataset, using a fixed subset of 5,000 data points, with 20% reserved for validation.
 To vary the dimensionality, an XGBoost model is fit to the original 388-dimensional dataset, and the most important features are retained according to XGBoost’s default feature selection criterion.
 The objective is to minimize the validation RMSE.
+In synthetic functions, additive Gaussian noise is introduced directly to the function evaluations (see Table 1 for details).
+The LCBench benchmarks rely on pre-trained surrogate models, where the posterior mean is evaluated, and in cases where $\sigma_{\varepsilon}$ is non-zero, additional noise is added to the evaluation of the posterior mean.
 Table 1: Noise levels for all Active Learning (AL) and Bayesian Optimization (BO) benchmarks.
@@
 Feature-reduced SVM BO 40 0.0
-In synthetic functions, additive Gaussian noise is introduced directly to the function evaluations.
-The LCBench benchmarks rely on pre-trained surrogate models, where the posterior mean is evaluated, and in cases where $\sigma_{\varepsilon}$ is non-zero, additional noise is added to the evaluation of the posterior mean.
-A.3 Compute Resources
-All experiments were conducted using an NVIDIA A40 GPU cluster.
-The compute usage to run all experiments in the main paper amounts to approximately 1000 GPU hours, and an additional 500 GPU hours to produce all the results provided in Appendix C.
-A.4 Licenses
+A.3 Licenses
 The following software packages, libraries and datasets were used in our experiments and for presenting the results in the paper:
@@
   • LCBench: Apache License
+A.4 Compute Resources
+All experiments were conducted using an NVIDIA A40 GPU cluster.
+The compute usage to run all experiments in the main paper amounts to approximately 1000 GPU hours, and an additional 500 GPU hours to produce all the results provided in Appendix C.
 Appendix B Derivation of HIPE as Joint Information Gain
@@
 These results provide additional insights into the robustness of the methods with respect to batch size during initialization and active learning, as well as their performance trade-offs across different evaluation metrics.
-C.1 Impact of Batch Size on Predictive Performance
+C.1 Runtime Analysis
+We compare acquisition function optimization runtime of HIPE and NIPV, and put that in relation to the runtime of the other components of a BO loop that employs fully Bayesian GP modeling.
+In Table 2 we demonstrate that while HIPE is substantially slower to optimize than NIPV, its total runtime still amounts to only a minor share of the total runtime of the overall BO run.
+Runtime Component (sec) HIPE NIPV Random
+Model fit (initial) 21.91 $\pm$ 0.15 21.91 $\pm$ 0.15 N/A
+Acquisition Opt. 19.61 $\pm$ 1.12 8.62 $\pm$ 0.37
+Model fit (BO loop) 181.79 $\pm$ 7.60 189.90 $\pm$ 5.17 177.33 $\pm$ 7.69
+qLogNEI Optimization 4.97 $\pm$ 0.18 4.44 $\pm$ 0.05 4.35 $\pm$ 0.06
+Table 2: Total runtime comparison (in seconds) for HIPE, NIPV, and Random across different components over the course of a 4-batch, $q=8$ optimization run.
+HIPE is more time-consuming to optimize compared to NIPV.
+While NIPV takes approximately half as long to optimize, the runtime to optimize HIPE is small relative to the time required to fit the fully Bayesian GP with NUTS.
+C.2 Impact of Batch Size on Predictive Performance
 We investigate how the initialization batch size $q$ affects predictive quality and inference performance on the LCBench tasks introduced in Section 5.3.
@@
 HIPE consistently leads in NLL and small-$q$ inference, while other methods achieve lower RMSE, indicating a trade-off between model calibration and pointwise prediction accuracy.
-C.2 Bayesian Optimization with Different Batch Sizes
+C.3 Bayesian Optimization with Different Batch Sizes
 We further investigate the impact of batch size on Bayesian optimization performance by evaluating all methods on the LCBench tasks with a reduced batch size of $q=8$.
@@
 HIPE is once again the best-performing method overall, but benefits from finding high-valued points during initialization to a greater extent than other methods.
-C.3 Active Learning with Different Batch Sizes
+C.4 Active Learning with Different Batch Sizes
 We analyze the effect of batch size in the active learning setting by comparing model performance under small batches ($q=8$) and large batches ($q=24$).
@@
 With larger batch sizes, HIPE exhibits improved relative performance, particularly on RMSE, demonstrating effective scalability under increased parallel evaluations.
-C.4 Complementary Hyperparameter Visualization on SVM
+C.5 Complementary Hyperparameter Visualization on SVM
 We display the inferred hyperparameter values of the left-out methods (Random, NIPV) on the 40D-SVM.